Assume the current interest rate on a one-year Treasury bond (1R1) is 2.17 percent, the current rate on a two-year Treasury bond (1R2) is 2.33 percent, and the current rate on a three-year Treasury bond (1R3) is 2.44 percent. If the unbiased expectations theory of the term structure of interest rates is correct, what is the one-year interest rate expected on T-bills during year 3 (E(3r1) or 3f1)?
Answer #1

E(3r1) = [(1 + 1R3)3 / (1 + 1R2)2] - 1

= [(1 + 0.0244)3 / (1 + 0.0233)2] - 1

= [1.0750 / 1.0471] - 1 = 1.0266 - 1 = 0.0266, or 2.66%